# Solving a Stackelberg Game on Transportation Networks in a Dynamic Crime Scenario: A Mixed Approach on Multi-Layer Networks

### v1

@@ 4.3 Theorem: Problem ”color covering” is $NP$-complete
    Output: ”yes” when there exists a path from the initial vertex with length $n$, such that all $m$ colors appear at least one of a vertex on the path. ”no” otherwise.
Theorem: Problem ”color covering” is NP-complete
@@
    Output: ”yes” if there exists an assignment of variables that makes all clauses true. ”no” otherwise.
@@
    If a clause with color $c$ contains $x_{i}$, the ”true” vertex for $x_{i}$ has color $c$.
@@
    If a clause with color $c$ does not contain $x_{i}$, the ”false” vertex for $x_{i}$ has color $c$.
@@ 5 Results and Discussion
In Fig. 2, we consider a sample network and use Dijkstra’s algorithm on the time-expanded network to generate the optimal attacker strategy. Here, we provide two mixed defender strategies as input with probabilities of 0.6 and 0.4. We demonstrate that our developed exact approach for the attacker can generate the optimal attacker strategy, enabling the attacker to escape without interdiction in a concise amount of time (see Table 1).
Figure 2: Optimal Attacker Strategy Design Using Dijkstra Algorithm on Time Expanded Network.
@@
Network Size: Crime Node: $0\_1$, Police Station: $0\_6$, $T_{max}$: 6, Exit Point: 5
@@
In Fig. 3, we consider a sample network and use the approximation algorithm on the time-expanded network to generate the near-optimal defender strategy. Here, we provide three pure attacker strategies as input with equal probability. We demonstrate that our developed approach for the defender can generate an efficient defender strategy that interdicts all attacker strategies quickly (see Table 2).
Figure 3: Defender Strategy Design Using Approximation algorithm.
@@
Network Size: Crime Node: $0\_1$, Police Station: $0\_6$, $T_{max}$: 6, Exit Point: 5
Test Case Final Defender strategy No. of Non-interdicted Attacker Strategies by the Final Defender Strategy Run Time (Sec)
@@
We compare the final defenders’ utility generated by the developed MLN-EIGS algorithm with the final defenders’ utility using MILP-EIGS algorithm, which uses the exact approaches named bestDo and bestAo, developed by Zhang et al. [3]. To establish a benchmark, we solve the Stackelberg Game using the bestDo and bestAo exact approaches. We compare both the solution quality and computational time.
@@
We show that for these small-size grids, our developed mixed approach named MLN-EIGS can provide an optimal solution like the MILP approach named MILP-EIGS in most cases and takes less computational time compared to MILP-EIGS.
6 Conclusion and Future Directions
In this paper, we consider a Stackelberg game theoretic model to solve security games considering the dynamic crime scenario. This paper proposes a novel optimal routing and resource allocation approach on a transportation network. The optimal strategy for the attacker is generated using an exact approach. To the best of our knowledge, this is the first time an exact approach using the multi-layer network concept is used for the attacker to solve a security game. This developed exact approach consists of Dijkstra’s algorithm and the layered graph concept. The near-optimal strategy for the defender is generated using a novel approximation algorithm and the layered graph concept on a transportation network in a time-efficient manner. The experiments are conducted on 3x3 to 9x9 grids.
The results are compared with the solutions developed using MILP approaches for both attacker and defender in this Stackelberg game scenario. For benchmarking, we formulate the Stackelberg game to provide optimal solutions using the adopted bestAO and bestDO approaches. The contribution of the paper is that the developed MLN-EIGS can offer quality solutions like MILP-EIGS but in a significantly shorter amount of time.
The developed approaches have a few limitations. For example, the proposed approach does not include real-time traffic. Future research can focus on developing efficient algorithms to incorporate traffic into the developed mixed approach. In addition, an exact approach for the defender can also be developed to generate an optimal strategy for the defender on a time-layered network.

### v2

@@ 4.3 Problem ‘color covering’ is $NP$-complete
    Output: ‘Yes’ if there exists a path from the initial vertex with length $n$, such that all $m$ colors appear on at least one vertex along the path. ‘No’ otherwise.
Theorem: Problem ‘color covering’ is NP-complete
@@
    Output: ‘Yes’ if there exists an assignment of variables that makes all clauses true. ‘No’ otherwise.
@@
    If a clause with color $c$ contains $x_{i}$, the ‘true’ vertex for $x_{i}$ has color $c$.
@@
    If a clause with color $c$ does not contain $x_{i}$, the ‘false’ vertex for $x_{i}$ has color $c$.
@@
5 MILP-EIGS benchmarking algorithm
To establish a benchmark, we formulate a novel Stackelberg game in which the defender acts as the leader and the attacker as the follower. The optimal strategies for both the attacker and defender are derived using bestAo and bestDo, which represent the MILP approaches for the attacker and defender, respectively (see Algorithm 3). These optimal approaches, developed by [3], provide the best strategies for attackers and defenders, given a predefined strategy set for each player. Similar to the MLN-EIGS approach, in this MILP-EIGS framework, the defender commits to the optimal mixed strategy. The attacker then uses the best defender mixed strategy as input to generate the optimal attacker strategy and updates his pure strategy set accordingly. The game utility is computed, and the process continues until the game utility remains below a threshold value $\epsilon$ for $K$ consecutive iterations. Once the game converges, the defender’s utility value is returned.
Output: $U_{d}:$ Defenders’ game utility;
repeat
       $(x,y)\leftarrow CoreLP(S^{\prime},A^{\prime})$;
       $BR:\;A^{*}\leftarrow{bestAo(x^{\star})}$;
       if $A^{*}\neq\emptyset$ then
             $A{{}^{\prime}}\leftarrow A^{{}^{\prime}}\cup A^{*}$;
       end if
      $S^{*}\leftarrow{bestDo(A^{{}^{\prime}})}$;
       if $S^{*}\neq\emptyset$ then
             $S^{{}^{\prime}}\leftarrow S^{{}^{\prime}}\cup S^{*}$;
       end if
until convergence;
return $U_{d}$.
Algorithm 3 MILP-EIGS benchmarking algorithm.
The vehicle interdiction problem is proved to be NP-hard ([3]). The best oracles, that is, the MILPs, encounter significant space and time complexity when applied to moderately large urban road networks with many nodes and edges. The MILP consists of bestDo for defenders and bestAo for the attacker. [3] develop these MILP approaches to find the best strategies.
5.1 MILP for attacker: bestAo
The MILP approach for the attacker (bestAo) constructs an optimal path for the attacker from the crime node to the exit node. The attacker’s utility decreases when more defender paths interdict this new attacker path. In this formulation, $A_{j,v}$ represents whether the attacker reaches node $v$ in the $j^{th}$ state. Equation (11) specifies that the attacker starts at node $v^{a}_{0}$, aims to reach the exit node $v_{\infty}$, and occupies only one node in each state. Equations (12) and (13) ensure that once the attacker reaches the sink node $v_{\infty}$, he remains there in subsequent states and confirms that the attacker’s strategy forms a valid path. Specifically, the attacker can only reach a neighboring node $u$ in state $j+1$ if they were at node $v$ in state $j$, as indicated by $\omega_{j,(v,u)}$ in Equations (14) and (15). Equations (16) and (17) initialize the time and update it in subsequent states. Finally, Equations (18) to (20) determine whether the attacker is intercepted by defender paths $S$, denoted by $z_{S}$.
$max\;\sum\limits_{s\in S^{{}^{\prime}}}(1-z_{s})x_{s}$ (10)
$s.t.\qquad A_{1,v^{a}_{0}}=1,A_{L^{a}_{max},v_{\infty}}=1,\sum\limits_{v\in V}A_{j,v}=1\qquad\forall j$ (11)
$A_{j+1,v_{\infty}}\geq A_{j,v_{\infty}}\qquad\forall j$ (12)
$\sum\limits_{u\in N(v)}A_{j+1,u}\geq A_{j,v}\qquad\forall v\in V,j$ (13)
$\omega_{j,(v,u)}\leq min(A_{j,v},A_{j+1,u})\qquad\forall(v,u)\in E,j$ (14)
$\omega_{j,(v,u)}\geq A_{j,v}+A_{j+1,u}-1\qquad\forall(v,u)\in E,j$ (15)
$t^{a}_{1}=0$ (16)
$t^{a}_{j+1}\geq t^{a}_{j}+\sum\limits_{(v,u)\in E}t_{(v,u)}\;\omega_{j,(v,u)}\qquad\forall j$ (17)
$-M\alpha^{j}_{S,r,i}\leq t^{r,in}_{i}-t^{a}_{j}\leq M(1-\alpha^{j}_{S,r,i})\qquad\forall S,r,i,j$ (18)
$-M\beta^{j}_{S,r,i}\leq t^{a}_{j}-t^{r,out}_{i}\leq M(1-\beta^{j}_{S,r,i})\qquad\forall S,r,i,j$ (19)
$z_{S}\geq\alpha^{j}_{S,r,i}+\beta^{j}_{S,r,i}+A_{j,v^{S^{r}}_{i}}-2\qquad\forall S,r,i,j$ (20)
$A_{j,v},\alpha^{j}_{S,r,i},\beta^{j}_{S,r,i},\omega_{j,(v,u)},z_{S}\in\{0,1\},t^{a}_{j}\in[0,t_{max}]$ (21)
5.2 MILP for defender: bestDo
The MILP approach for defenders (bestDo) is described below, focusing on maximizing the rate at which the defender can intercept the attacker within a specified time frame in a large transportation network. The bestDo provides the optimal formulation for the defender’s movements over time. It devises a path for the defender that intercepts the maximum number of attacker paths, thereby maximizing the defender’s utility. In this approach, the attacker attempts to flee the city by utilizing any exit node $v_{\infty}$ from the crime scene, represented as node $v^{0}_{a}\in V$. A defender’s state is defined by the tuple $(v^{r},t^{r,in},t^{r,out})$, indicating the defender’s presence at node $v^{r}$ during the time interval $[t^{r,in},t^{r,out}]$. The waiting time at a node is calculated as $t^{r,out}-t^{r,in}=k\delta$, and the travel time between consecutive nodes is estimated as $t^{r,in}_{i+1}-t^{r,out}_{i}=dist(v^{r}_{i},v^{r}{i+1})/sp_{d_{r}}$.
In this context, $s^{r}_{i,v}=1$ signifies that defender $d_{r}$ reaches node $v$ at the $i^{th}$ state of $S^{r}$. Equation (23) specifies that the initial position of $d_{r}$ is $v^{r}_{0}$, and $d_{r}$ remains on a single node at each state of $S^{r}$. In Equations (24) and (25), the $0/1$ variable $\omega_{r,i,(v,u)}$ indicates whether a path exists between nodes $v$ and $u$ between the $i^{th}$ and $(i+1)^{th}$ states of $S^{r}$. Equation (26) specifies that the strategy begins at time 0 and ends at $t_{max}$, where $L_{max}$ represents the maximum allowable length of the defender’s strategy sequence, and the time the defender remains at the $i^{th}$ state is $k_{r,i}\delta$. Equation (27) guarantees that the defender travels between stops along the shortest path. With Eqs.(28)-(32), $z_{A}$ indicates whether the attacker using path $A$ is intercepted.
Specifically, $(v^{A}_{j},t^{A}_{j})$ represents the attacker’s $j^{th}$ state in $A$, while $\gamma^{A,j}_{r,i}$ indicates whether the attacker using path $A$ encounters defender $d_{r}$ at the attacker’s $j^{th}$ position while $d_{r}$ is at their $i^{th}$ stop. Additionally, $\alpha^{A,j}_{r,i}$ and $\beta^{A,j}_{r,i}$ denote whether the attacker arrives at their $j^{th}$ position after $d_{r}$ arrives or before $d_{r}$ leaves, respectively. The constant $M$ is a large number used in the formulation.
$max\;-\sum\limits_{A\in A^{{}^{\prime}}}(1-z_{A})y_{A}$ (22)
$s.t.\qquad s^{r}_{1,v^{r}_{0}}=1,\sum\limits_{v\in V\setminus\{v_{\infty}\}}s^{r}_{i,v}=1\qquad\forall r,i$ (23)
$\omega_{r,i,(v,u)}\leq min(s^{r}_{i,v},s^{r}_{i+1,u})\qquad\forall r,i,u,v$ (24)
$\omega_{r,i,(v,u)}\geq s^{r}_{i,v}+s^{r}_{i+1,u}-1\qquad\forall r,i,u,v$ (25)
$t^{r,in}_{1}=0,t^{r,out}_{L^{d}_{max}}=t_{max},t^{r,out}_{i}=t^{r,in}_{i}+k_{r,i}\delta\qquad\forall r,i$ (26)
$t^{r,in}_{i+1}=t^{r,out}_{i}+\sum\limits_{v,u\in V\setminus\{v_{\infty}\}}dist(v,u)\omega_{r,i,(v,u)}\qquad\forall r,i$ (27)
$-M\alpha^{A,j}_{r,i}\leq t^{r,in}_{i}-t^{A}_{j}\leq M(1-\alpha^{A,j}_{r,i})\qquad\forall r,i,A,j$ (28)
$-M\beta^{A,j}_{r,i}\leq t^{A}_{j}-t^{r,out}_{i}\leq M(1-\beta^{A,j}_{r,i})\qquad\forall r,i,A,j$ (29)
$\gamma^{A,j}_{r,i}\leq(\alpha^{A,j}_{r,i}+\beta^{A,j}_{r,i}+s^{r}_{i,v^{A}_{j}})/3\qquad\forall r,i,A,j$ (30)
$\gamma^{A,j}_{r,i}\geq\alpha^{A,j}_{r,i}+\beta^{A,j}_{r,i}+s^{r}_{i,v^{A}_{j}}-2\qquad\forall r,i,A,j$ (31)
$z_{A}\leq\sum\limits_{j,r,i}\gamma^{A,j}_{r,i}\qquad\forall A$ (32)
$s^{r}_{i,v},\omega_{r,i,(v,u)},\alpha^{A,j}_{r,i},\beta^{A,j}_{r,i},\gamma^{A,j}_{r,i},z_{A}\in\{0,1\}$ (33)
$k_{r,i}\in Z_{\geq 0},t^{r,in}_{i},t^{r,out}_{i}\in[0,t_{max}]$ (34)
To assess optimality, we compare the utility and computational time of the developed MLN-EIGS with that of MILP-EIGS.
@@ 6 Results and discussion
In Fig. 2, we consider a sample network of 4 nodes in which police stations are nodes 2 and 3, the crime node is 1, the maximum time limit ($t_{max}$) is 5, and the exit point is node 4. Here, $0\_4$ indicates node 4 at timestamp 0 $(t=0)$ in the multi-layer network. In this example, we provide two mixed defender strategies as input with probabilities of $1/3$ and $2/3$. Each node in this multi-layer network is assigned a corresponding probability of interdiction. To generate the optimal attacker strategy, we use Dijkstra’s algorithm on the time-expanded network (see Fig. 3). The final attacker strategy is represented by the red line in the multi-layer network, which follows the path $0\_1\rightarrow 3\_3\rightarrow 5\_4$. We demonstrate that our developed exact approach for the attacker can generate the optimal attacker strategy, enabling the attacker to escape without interdiction in a concise amount of time (see Table 1).
Figure 2: Sample network for designing the optimal attacker strategy.
Figure 3: Design of a multi-layer network for attacker considering a sample network (Fig. 2).
@@
Network Size: 4 Nodes, Crime Node: $0\_1$, Police Stations: $0\_2$, $0\_3$, $T_{max}$: 5, Exit Point: 4
@@
In Fig. 4, we consider a sample network of 6 nodes in which the police station is node 6, the crime node is 1, the maximum time limit $t_{max}$ is 6, and the exit point is node 5. Here, $0\_6$ indicates node 6 at timestamp 0 $(t=0)$ in the multi-layer network. We input three attacker strategies, each with an equal probability. Each node within the same attacker strategy is colored identically in this multi-layer network. We use an approximation algorithm on the time-expanded network to generate the near-optimal defender strategy (see Fig. 5). The final defender strategy is represented by the green curvy lines in the multi-layer network, which follows the path $0\_6\rightarrow 2\_3\rightarrow 4\_4\rightarrow 5\_5\rightarrow 6\_5$ and $0\_6\rightarrow 2\_3\rightarrow 4\_5\rightarrow 5\_5\rightarrow 6\_5$. We demonstrate that our developed approach for the defender can generate an efficient defender strategy that interdicts all attacker strategies quickly (see Table 2).
Figure 4: Sample network for designing near-optimal defender strategy.
@@
Network Size: 6 Nodes, Crime Node: $0\_1$, Police Station: $0\_6$, $T_{max}$: 6, Exit Point: 5
Test Case Final Defender strategy Utility of the Final Defender Strategy Run Time (Sec)
@@
Figure 5: Design of a multi-layer network for defender considering a sample network (Fig. 4).
We compare the final defenders’ utility generated by the developed MLN-EIGS algorithm with the final defenders’ utility using the MILP-EIGS algorithm, which employs the exact approaches named bestDo and bestAo, developed by [3]. To establish a benchmark, we solve the Stackelberg Game using the bestDo and bestAo MILP approaches. In both cases, the Stackelberg Game converges after $K$ consecutive iterations where the game utility is below $\epsilon$. Here, the values of $K$ and $\epsilon$ are assigned to 10 and 0.05, respectively. We compare both the solution quality and computational time.
@@
In the above tables, we consider seven grids ranging in size from 3*3 to 9*9 and evaluate the utility and runtime for the MLN-EIGS and MILP-EIGS approaches by varying the crime node location, initial defender positions, $t_{max}$ (maximum runtime), and exit node. In the case of a 3*3 grid, MLN-EIGS takes slightly less computational time with a performance gap of zero, whereas for the 4*4 grid, MILP-EIGS takes nearly 10 seconds longer than MLN-EIGS with no performance gap. It is evident that MLN-EIGS provides optimal results similar to MILP-EIGS for smaller grids. In the case of the 5*5 grid, MILP-EIGS takes more than 20 seconds longer, and for the 6*6 grid, the time difference exceeds 30 seconds, sometimes reaching 70 seconds. MLN-EIGS performs as well as MILP-EIGS for both the 5*5 and 6*6 grids, with only one test case showing a performance gap. For the 7*7 and 8*8 grids, while the performance gap remains zero, there is a significant difference in computational time between the two approaches. On the 7*7 grid, MILP-EIGS takes over one minute, and on the 8*8 grid, it takes over five minutes longer than MLN-EIGS. For the 9*9 grid, although test case four shows a performance gap of 0.5, MLN-EIGS still outperforms MILP-EIGS in terms of computational efficiency. Overall, the results indicate that MLN-EIGS consistently provides a time-efficient, high-quality solution, while MILP-EIGS struggles to offer a time-efficient solution for the developed Stackelberg game.
7 Conclusion
In this paper, we present a Stackelberg game model to address security challenges in a dynamic crime scenario. We propose a novel optimal routing and resource allocation approach on a transportation network. The attacker’s optimal strategy is determined using an exact method, which, to our knowledge, is the first to apply the multi-layer network concept in solving a security game. This exact approach combines Dijkstra’s algorithm with the layered graph model. The defenders’ near-optimal strategy is generated through a novel approximation algorithm, utilizing the layered graph concept on a transportation network in a time-efficient way.
We solve the developed Stackelberg game using the adopted MILP approaches to provide optimal solutions. For benchmarking, we compare the results of our approach with those generated using MILP methods. The key contribution of this work is that our MLN-EIGS approach delivers high-quality solutions comparable to MILP-EIGS but in significantly less time.
There are some limitations, such as the exclusion of real-time traffic. Future research could focus on developing efficient algorithms to incorporate traffic data and on creating an exact approach for the defender to generate optimal strategies using the time-layered network concept.
Acknowledgements We are grateful to the members of the Multi-Agent Laboratory at Kyushu University for their insightful discussions and comments. This research is funded by a project supported by the Grants-in-Aid for Scientific Research from the Japan Society for the Promotion of Science.